OpenStudy (anonymous):

A sheet of metal is 32.7mm wide and 80.4mm long. If it weighs 5.501g and the density of the metal is 2.70g/cc, what is the thickness of the sheet? (in mm) I know that the density formula is d=m/V and that to get the thickness of the sheet I would need to find the height which I can use the volume formula lxwxh, but I am unsure how to convert cc or to get rid of it i just don't know

OpenStudy (anonymous):

Okay I did this in chemistry. The lab was called thickness of a thin layer. so basically we know two sides that gives us area. which is two dimensions and we need the other one for three dimensions. A of the sheet is 32.7 mm * 80.4 mm= 2629.08 Now we need to find the volume from the density given and the mass but the problem is we have cc and mm so we have to convert it first 1 cc = 1000mm^3 2.70/1000 new line is 0.0027g/mm^3 now we can get the volume from the mass 5.501g x 1mm^3/0.0027g grams (we wants the volume so V = m/d) 2037.41 mm^3 is what we get now to get the layer lets remove two dimensions. Which was the area. to get the unkown dimension (height) 2037.41mm^3/ 2629.08 mm^2 Height = 0.77 mm

OpenStudy (anonymous):

if 1cc is = to 1000mm^3 then why would 2.70g/cm^3 be 0.0027g/mm^3? Wouldn't it be 2700mm^3?

OpenStudy (anonymous):

now all we did was change the unit on the bottom. so 1 cc= 1000mm i divided it because it was under the 2.70g bar. how do you have 2700mm^3? you can't just get rid of the g. then it wouldn't be a density.

OpenStudy (anonymous):

i meant to put the g sorry, but what do you mean by under the 2.70g bar? If it's 2.70g/cc 1cc=1000mm^3 still. I don't understand that

OpenStudy (anonymous):

2.70 g/cc is a fraction under the fraction bar